Circular No. 7418

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 J2 (SOHO) D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of another comet found in LASCO C2 data by M. Oates via the SOHO website. The comet brightened from V = 7.8 on May 7.254 UT to 7.1 on May 7.368, but had faded to V = 8.3 on May 7.452, and it was not seen in the C3 instrument. No tail was evident. Astrometry (measured by Biesecker, reduced by B. G. Marsden) and orbital elements appear on MPEC 2000-J32; this comet is evidently not a Kreutz sungrazer.

2000 UT  α_{2000}  δ_{2000}
May 7.243  3^{h}01.^{m}7  +17^{o}22′

XTE J1118+480 M. Uemura, Kyoto University, reports on behalf of the VSNET XTE

J1118+480 collaboration team: “Unfiltered CCD photometry of this xray transient (cf. IAUC 7389) has been undertaken since Mar. 30 by numerous observers around the world (M. Uemura, T. Kato, K. Matsumoto, M. Honkawa, L. Cook, B. Martin, G. Masi, A. Oksanen, M. Moilanen, R. Novak, and Y. Sano). Analysis of the combined light curve (covering 37 days) revealed the growth of the periodic variation profile from sinusoidal to ‘asymmetric humplike’. The period has significantly changed from 0.1709(3) day during Mar. 30.7–Apr. 1.4 UT to 0.1703(1) day during Apr. 26.8–May 5.7. The period change strongly suggests that the optical modulation not only reflects the orbital modulation but also shows for the first time ongoing development of superhumps in a blackhole transient (e.g., Bailyn 1992, Ap.J. 391, 298; Mineshige et al. 1992, PASJ 44, L15; O’Donoghue and Charles 1996, MNRAS 282, 191). We invite worldwide observers to join the VSNET collaboration team as part of a simultaneous multiwavelength campaign. An ASCA x-ray pointing observation (led by Y. Ueda) in the band 0.5–10 keV is planned between May 11.5 and 12.0. More detailed information can be seen at http://www.kusastro.kyoto-

u.ac.jp/vsnet/Xray/xtej1118-camp.html.”

COMET C/1999 S4 (LINEAR) This comet was detected low in the morning sky by K. Kadota, Ageo,

Saitama, Japan (0.18-m reflector + CCD) at m_{1} ∼ 13.0 on May 4.78 UT (coma diameter 0′.4).

2000 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green